OpenStudy (anonymous):

convert to exponential form: log3 81=4 => a^b=c

OpenStudy (anonymous):

\[\log_{3} 81=4 => a^b=c\]

jhonyy9 (jhonyy9):

3 on exponent 4 is 81

jhonyy9 (jhonyy9):

than there will be 4log3(3)=4

OpenStudy (anonymous):

so is that the answer?

OpenStudy (anonymous):

\[3^4=81\]

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

how did you get that?

jhonyy9 (jhonyy9):

3*3*3*3=81

OpenStudy (anonymous):

mine is correct

OpenStudy (anonymous):

ohhhh i see

OpenStudy (anonymous):

\[\log_a c=b\] every logarithm can be rewritten like this: \[a^b=c\]

OpenStudy (anonymous):

tomas: u are correct but idk how to get that ... is it wat jhonyy said?

OpenStudy (anonymous):

don't ever forget \[b^x=y\iff log_b(y)=x\]